Exercise 3. Let α and β be two real numbers so that α ≤ β and β ≤ α. Then prove α = β.


1
Expert's answer
2021-01-12T02:33:18-0500

"\alpha \leqslant \beta" "\implies""\alpha \isin (-\infty,\beta]"

"\beta \leqslant \alpha" "\implies" "\alpha\isin [\beta, \infty)"

"\implies \alpha \isin (-\infty,\beta] \bigcap [\beta,\infty)=[\beta\beta]=\{\beta\}"

"\implies \alpha=\beta"